OpenStudy (anonymous):

1. Find the actual zero of f(x) = x3 + 5x2 - 4x - 20? i really need help on this :(

OpenStudy (loser66):

Any idea? Hint: have to use "try and error" method. plug \(\pm1, \pm2\) to get the first zero, then use long division or synthetic to get the leftover.

OpenStudy (anonymous):

i have no idea what any of that is.

OpenStudy (loser66):

Show me your trying, please, plug x = 1, x =-1, x =2, x=-2 into the original one, what do you get for f(x)?

OpenStudy (anonymous):

You can rearrange the terms and do some factoring: x^3 + 5x^2 - 4x - 20 =x^3-4x+5x^2-20 =x(x^2-4)+5x^2-20 =x(x^2-4)+5(x^2-4) =(x+5)(x^2-4) =(x+5)(x+2)(x-2) So zeroes are: -5,-2,+2